use the laplace transform method to solve the boundary value problem

∂u/∂t=3(∂u/∂x), u(x,0)=4e-2x

1
Expert's answer
2021-10-21T10:18:25-0400

Appliing Laplase transworm with respect to t we have

U(x,p)p-u(x,0)="3\cdot U(x,p)'_x"

or

"\frac{d}{dx}U(x,p)=U(x,p)\cdot \frac{p}{3}-\frac{4}{3}\cdot e^{-2x}"

homogeneous equation has the general solution

"U(x,p)=Ce^\frac{px}{3}"

Let C=C(x).

Then

"C'(x)e^\frac{px}{3}=-\frac{4}{3}\cdot e^{-2x};"

C'(x)="-\frac{4}{3}\cdot e^{(-2-\frac{p}{3})x}"

"C(x)=\frac{4}{6+p}e^{(-2-\frac{p}{3})x}"

U(x,p)="\frac{4}{6+p}e^{-2x}"

Using table of Laplase transform we have

"\frac{4}{p+6}\doteqdot 4\cdot e^{-6t}"

Finalyy we have

u(x,t)="4\cdot e^{-(6t+2x)}"